OpenStudy (anonymous):

x/5 - x/6 =2 Find X.

OpenStudy (anonymous):

Bring the x to a common divisor, then solve.

OpenStudy (homeworksucks):

Multiply everything by a number which is a multiple of both 5 and 6, and solve from there

OpenStudy (anonymous):

good, easy solution, yes

OpenStudy (anonymous):

im honestly lost

OpenStudy (anonymous):

ca you guys show me work, so i can get a visual

OpenStudy (anonymous):

\[\frac{x}{5} - \frac{x}{6} = 2\] if you stick to Homeworksucks' solution, multiply everything with a common multiple of 5 and 6, let's take 30: \[\frac{x*30}{5} - \frac{x*30}{6} = 2*30 \] Now simplify and solve!

OpenStudy (anonymous):

is the answer 165

OpenStudy (anonymous):

How did you get to this one?

OpenStudy (anonymous):

i cross multiplied

OpenStudy (anonymous):

way too complicated: \[\frac{30x}{5} = 6x\] \[\frac{30x}{6} = 5x\] So \[6x-5x = 1x = x = 2*30 = 60\]

OpenStudy (anonymous):

i understood until the 2*30=60

OpenStudy (anonymous):

Well that's multiplication?

OpenStudy (anonymous):

i know but why did you do it

OpenStudy (anonymous):

you need x, and i dont understand why you did that for x

OpenStudy (anonymous):

To keep the equation correct you have to multiply EVERYTHING with 30, not just the terms with x

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

x/4 - x/12 =20 is X 120 then

OpenStudy (anonymous):

yes, you got it

OpenStudy (anonymous):

omg thank you!

OpenStudy (anonymous):

To validate your results, just plug in your solution, in this case: 120/4 - 120/12 = 20-10 = 20

OpenStudy (anonymous):

yeah thank you

OpenStudy (anonymous):

You're very welcome.